OpenStudy (annalee607):

Use Euler's Formula to find the missing number. Vertices:13 Edges:2 Faces:?

OpenStudy (jpharper12345678cars):

is edges supose to be 28?

OpenStudy (jpharper12345678cars):

if it is i have that answer

OpenStudy (annalee607):

yes

OpenStudy (jpharper12345678cars):

its 17

OpenStudy (jpharper12345678cars):

r = m - v + 2, where r = faces, v = vertices, and m = edges r = 28 - 13 + 2 r = 15 + 2 r = 17,

OpenStudy (jpharper12345678cars):

do you have anymore you need help with?

OpenStudy (jpharper12345678cars):

tog me if you do